# TRABSA: Interpretable Sentiment Analysis of Tweets using Attention-based BiLSTM and Twitter-RoBERTa

### v1

@@ Abstract
Sentiment analysis is a pivotal tool in understanding public opinion, consumer behavior, and social trends, underpinning applications ranging from market research to political analysis. However, existing sentiment analysis models frequently encounter challenges related to linguistic diversity, model generalizability, explainability, and limited availability of labeled datasets. To address these shortcomings, we propose the TRABSA model, a novel hybrid sentiment analysis framework that integrates transformer-based architectures, attention mechanisms, and BiLSTM networks. Leveraging the latest RoBERTa-based transformer model trained on a vast corpus of 124M tweets, our research bridges existing gaps in sentiment analysis benchmarks, ensuring state-of-the-art accuracy and relevance. Furthermore, we contribute to data diversity by augmenting existing datasets with 411,885 tweets from 32 English-speaking countries and 7,500 tweets from various US states. This study also compares six word-embedding techniques, identifying the most robust preprocessing and embedding methodologies crucial for accurate sentiment analysis and model performance. We meticulously label tweets using three distinct lexicon-based approaches and select the best one, ensuring optimal sentiment analysis outcomes and model efficacy. Here, we demonstrate that the TRABSA model outperforms the current seven traditional ML models, four stacking models, and four hybrid deep learning models, yielding notable gain in accuracy (94%) and effectiveness with a macro average precision of 94%, recall of 93%, and F1-score of 94%. Our further evaluation involves two extended and four external datasets, demonstrating the model’s consistent superiority, robustness, and generalizability across diverse contexts and datasets. Finally, by doing a thorough study with SHAP and LIME approaches, we offer insights into the interpretability of the TRABSA model, improving comprehension and confidence in the model’s predictions. Our study results make it easier to analyze how citizens respond to resources and events during pandemics since they are integrated into a decision-support system. Applications of this system provide essential assistance for efficient pandemic management, such as resource planning, crowd control, policy formation, vaccination tactics, and quick reaction programs.
@@
Sentiment analysis , Lexicon-based labeling , Attention mechanism , eXplainable AI (XAI) , SHAP , LIME
journal: Expert Systems with Applications
1 Introduction
@@
Due to the growth of textual data on social media platforms, news stories, reviews, and consumer feedback, sentiment analysis (SA), a crucial aspect of natural language processing (NLP), has seen growing attention and usage across several domains [Zhang & Liu, 2017]. Institutes may get crucial insights into public opinion, consumer preferences, market trends, and brand impression by identifying and analyzing feelings conveyed in text [Chaturvedi et al., 2017]. Thus, SA plays a critical role in directing marketing initiatives, product development, company strategies, and reputation management [Taboada, 2016]. Additionally, SA is useful in a variety of domains, including politics, healthcare, economics, and the social sciences, where decision-making and policy development depend on a knowledge of human emotions and attitudes.
@@
Even with its broad use, SA still has issues that need for more study and creativity. The lack of generalizability and robustness of SA models is one of the primary issues, especially when applying them to a variety of languages, domains, and datasets [Kallam et al., 2023]. Because existing models frequently display different performance levels based on the properties of the data, they are less dependable in real-world situations where the data distribution may change greatly [Kumar, 2022]. Furthermore, SA models’ interpretability is still a major worry, particularly in high-stakes scenarios when model predictions are used to make judgments [Jawale & Sawarkar, 2020]. Deep learning (DL) models are black-box in nature, which makes it difficult to grasp how these models get to their conclusions. This makes it difficult to implement, impedes trust, and holds people accountable for important decision-making processes.
@@
Inspired by these difficulties, the goal of this study is to develop a strong, broadly applicable, and easily interpreted model of SA that will overcome the shortcomings of current techniques. Through utilizing improvements in DL, attention mechanisms, and interpretability methodologies, our goal is to develop a model that performs well on a variety of datasets and offers insights into how it makes decisions. This research advances the area of SA by bridging the gap between model performance, interpretability, and practical application. Our aim is to improve the trustworthiness, transparency, and usefulness of SA models using empirical assessments and interpretability studies. This will enable enterprises to make well-informed decisions by relying on dependable sentiment insights.
@@
Our research aims to fill several critical gaps in the existing literature. First off, although SA has received a lot of attention—especially when it comes to social media data—more robust and interpretable models are still required in order to classify sentiments across various languages and domains accurately. Many current methods are not transparent, scalable, or generalizable, making it difficult to use them in real-world situations. Furthermore, a notable deficiency in current datasets for SA is the absence of representation for various English language usage patterns. Variations in vocabulary, grammar, and contextual usage of English across national boundaries result in subtle discrepancies in the presentation of distinct emotions. This variability presents a problem for SA models, which can find it difficult to assess sentiments in various language circumstances appropriately. More advanced methods are required to capture minute semantic subtleties and adjust to changing contextual signals since current models may not be able to comprehend sarcasm, context-dependent sentiment changes, or nuanced sentiment expressions.
@@
This study addresses the need for robust and generalizable SA models by proposing the “Transformer-based Attention Bidirectional LSTM for Sentiment Analysis (TRABSA)” model. The TRABSA model integrates the strengths of transformer-based architectures with attention mechanisms and bidirectional long short-term memory (BiLSTM) networks to enhance the performance and adaptability of SA tasks. Our method seeks to capture the broad diversity of English language usage and offer a more thorough knowledge of sentiment expression in various linguistic situations by combining data from several locations into a single dataset. By using this method, TRABSA can more effectively adjust to the subtle differences in English language usage across various groups, which improves the precision and significance of SA findings. We test the performance of the TRABSA model on a range of DL architectures and datasets, including extensive Twitter and external social media datasets, with a particular emphasis on scalability, accuracy, and consistency. Additionally, we do interpretability assessments utilizing the SHAP and LIME approaches to understand the model’s decision-making mechanism. We show the TRABSA model’s generalizability and robustness through our thorough examination, providing a viable method for SA in various real-world situations.
Our research makes nine-fold key contributions:
@@
    We propose the TRABSA model, a novel hybrid SA framework that combines transformer-based architectures, attention mechanisms, and BiLSTM networks to improve SA performance.
@@
    This research leverages the latest RoBERTa-based transformer model, trained on a vast corpus of 124M tweets, to bridge existing gaps in SA benchmarks, ensuring state-of-the-art accuracy and relevance.
@@
    We extended the existing dataset by scraping 411,885 tweets from 32 English-speaking countries to include diversity in the Global Twitter COVID-19 Dataset, acknowledging the varied perspectives and discourse across regions.
@@
    We scraped an additional 7,500 tweets from different states of the USA to deepen geographical representation in the USA Twitter COVID-19 Dataset, allowing for localized insights and analysis.
@@
    This article thoroughly compares word embedding techniques, establishing the most robust preprocessing and embedding methodologies essential for accurate SA and model performance.
@@
    We methodically label tweets using three distinct lexicon-based approaches and rigorously select the most effective one, ensuring optimal SA outcomes and model efficacy.
@@
    We conduct extensive experiments to assess the TRABSA model’s performance on the UK Twitter Dataset, benchmarking against 7 traditional ML models, 4 stacking models, and 4 DL models, demonstrating its superiority and versatility.
@@
The rest of this article is structured as follows: Section “Related Works” reviews state-of-the-art literature in SA using ML-DL and interpretability techniques. In section “Methodology”, the data collection and preprocessing techniques, unsupervised text labeling, implemented ML and DL models for benchmarking, architecture, and methodology of the proposed TRABSA model are described. The section “Experimental Analysis and Results” presents the experimental setup, evaluation metrics, results, and robustness analysis of the TRABSA model. The SHAP and LIME analysis conducted on the TRABSA model are covered in section “Interpretability Analysis”. Section “Disucssions” addresses the findings and implications of our investigation. In conclusion, the article is summarized, and future research directions are outlined in section “Conclusions and Future Directions”.
2 Related Works
@@
When it comes to classifying data into positive, neutral, and negative sentiment polarity, SA is essential. Exploring a wide range of emotions is the focus of the emerging domains of SA [Cambria, 2016]. Sentiments can be further classified into categories like satisfaction and rage within certain settings, such as political disputes [D’Andrea et al., 2015]. The development of SA approaches with ambivalence management has allowed for the classification of sentiments into distinct classes, including sorrow, anger, anxiety, excitement, and happiness, leading to more nuanced outcomes [Wang et al., 2020]. While SA has typically focused on textual data, it has expanded to include multimodal SA, which explains data from devices that employ audio- or audio-visual formats [Yang et al., 2022]. The extension of SA into multimodal analysis highlights its variety and complexity, creating opportunities for a wide range of NLP applications. The variety of options is further highlighted by the fast growth of NLP, which is fueled by research in neural networks [Ray & Chakrabarti, 2022]. Notably, the development of Neurosymbolic AI, which combines symbolic reasoning and DL, offers a viable method of improving NLP capabilities [Sarker et al., 2022], highlighting the various paths that NLP research is taking. Lexicon-based methods, machine-learning (ML)–based methods, and hybrid techniques are the three main methodologies used to solve text categorization and emotion detection challenges. Word polarity is used by lexicon-based approaches, and machine learning (ML) techniques see text analysis as a classification problem that may be further divided into supervised, semi-supervised, and unsupervised learning approaches [Aqlan et al., 2019]. In real-world applications, SA results are frequently improved by combining ML with lexicon-based techniques.
@@
In Ahmed & Ahmed’s work, positive and negative emotions were used to classify gathered fake newspapers using a variety of approaches, including TF-IDF, random forest (RF), Naïve Bayes (NB), etc. [Ahmed & Ahmed, 2023]. According to their results, out of all the classifiers used, the Naïve Bayes classifier had the best accuracy (89.30%). In order to identify feelings in the Twitter sentiment 140 datasets, Gaur et al. used TF-IDF feature extraction and the Naïve Bayes Classifier [Gaur et al., 2023]. The model produced improved accuracy (84.44%) and precision when measured using several performance criteria, such as accuracy, recall, and precision. The COVID-19-related data that Qi & Shabrina (2023) examined came from Twitter users in major English cities [Qi & Shabrina, 2023]. They conducted a comparative analysis of ML models, including Vader and Textblob, RF, support vector classification (SVC), and multinominal Naïve Bayes (MNB) models. According to the results of their investigation, SVC with TF-IDF demonstrated better accuracy than the other models. In order to assess opinions about Saudi cruises, Al Sari et al. created three different datasets from social media platforms [Al sari et al., 2022]. With oversampled Snapchat data, they used ML techniques, including RF, MLP, NB, voting, SVM, and the n-grams feature extraction approach to reach 100% accuracy with the RF algorithm. A customized approach for explicit negation detection was presented by Mukherjee et al. [Mukherjee et al., 2021]. They used TF-IDF for feature extraction and a variety of ML techniques, including as NB, SVM, and Artificial Neural Networks (ANN), to analyze sentiment in Amazon reviews. According to their research, ANNs using negative classifiers had the best accuracy (96.32%). Using reviews from an international hotel, Noori developed a unique algorithm for classifying client sentiment [Noori, 2021]. Following the processing of the reviews, document vectors were created using the TF-IDF extractor and trained using SVM, ANN, NB, k-nearest neighbor (KNN), decision tree (DT), and C4.5 models. Outperforming other models, the DT model scored the highest accuracy (98.9%) with 1800 features. Using N-gram extraction, Zahoor & Rohilla compared NB, SVM, RF, and long short-term memory networks (LSTM) classifiers on preprocessed datasets [Zahoor & Rohilla, 2020]. In the majority of datasets, including the BJP and ML Khattar datasets, NB showed the best accuracy. In order to turn COVID-19-related tweets into a text corpus and determine the most common terms using N-grams, Samuel et al. used logistic regression (LR) and NB models [Samuel et al., 2020]. Their results showed that for short tweets, NB and LR had peak accuracy rates of 91% and 74%, respectively. For lengthier tweets, both models performed pretty poorly. Using Maximum Entropy (ME), SVM, and LSTM models, Kumar et al. examined the effects of age and gender on customer reviews [Kumar et al., 2020]. LSTM used word2vec, but the NB, ME, and SVM algorithms used Bag of Words (BOW) feature extraction. For female data, the over-50 age group showed the highest accuracy. SVM and MNB with TF-IDF extraction were used by Zarisfi Kermani et al. on four Twitter datasets, and they proposed semantic scoring techniques to represent features in the vector space [Zarisfi Kermani et al., 2020]. According to their findings, the suggested technique outperformed the MNB algorithm in three datasets, with the STS dataset showing the greatest MNB performance.
@@
In 13 languages with different Indic scripts, Bansal et al. looked at the identification of objectionable language [Bansal et al., 2022]. They assessed four sophisticated transformer-based models and contrasted the Transformer-based method with traditional ML models. Out of all of them, XLM-RoBERTa with BiGRU performed better. Furthermore, adding emoji embeddings to XLM-RoBERTa improved the model’s efficacy even further. Due to the combined dataset’s code-mixing, training using datasets from 13 Indic languages performed better than training with separate models. Gupta et al. [Gupta et al., 2021a] presented a unique emotion analysis approach for real-time COVID-19 tweets, examining eight emotions in different domains. The analysis of tweets from India showed changes in emotional reactions, such as less happiness except for nature. Because of their commitment, teachers’ faith in education has grown. In terms of precision and recall, the method put out by Gupta et al. produced aspect-based graphical and textual summaries from mobile reviews [Gupta et al., 2019], outperforming baseline approaches. Using Twitter data from the Delhi Election 2020, Gupta et al. conducted political echo chamber experiments and investigated the elements that contribute to the creation of echo chambers as well as the role played by users of opposing parties in promoting partisan material [Gupta et al., 2022]. Gupta and colleagues employed ML algorithms and lexicon-based methods to assess sentiment in Hindi tweets. They found that an integrated CNN-RNN-LSTM model produced an accuracy of 85% [Gupta et al., 2021b]. Basiri et al. investigated attitudes about the epidemic in eight different nations using DL classification algorithms, and their findings showed distinct sentiment patterns and relationships with pandemic indicators [Basiri et al., 2021]. Using the BERT model, Hayawi et al. achieved excellent accuracy in their ML-based method for spotting COVID-19 vaccination disinformation [Hayawi et al., 2022]. Using BERT, Vishwamitra et al. were able to detect hate speech connected to elderly individuals and the Asian community on Twitter during the pandemic [Vishwamitra et al., 2020]. They were able to identify separate word connections for various hate speech datasets. Before and after the initial COVID-19 case announcement, Chen et al. monitored conversations in Luxembourg about policy and daily life; post-announcement, travel-related issues dominated, perhaps because of the region’s large immigrant population [Chen et al., 2021]. In order to emphasize changing emotions over time, Kabir et al. used ML for word extraction and emotion categorization in COVID-19 tweets [Kabir & Madria, 2021]. In order to categorize COVID-19-related Twitter postings, Valdes et al. created a BERT-based model [Valdes et al., 2021], proving the use of domain-specific data for improved performance. During the pandemic, Tziafas et al. used an ensemble architecture to recognize false information, using transformer-based encoders to achieve high accuracy [Tziafas et al., 2021]. High assessment scores were obtained by Sadia et al. utilizing BERT to conduct SA of COVID-19 tweets [Sadia & Basak, 2021]. Song et al. analyzed several facets of misinformation diffusion and created a model to categorize misinformation related to COVID-19 [Song et al., 2021]. Using NLP models, Hossain et al. assessed a dataset for COVID-19-related misinformation detection, offering preliminary benchmarks for advancement [Hossain et al., 2020]. During the COVID-19 lockdown, Chintalapudi et al. used BERT to assess sentiment in Indian tweets, and they showed better accuracy than other models [Chintalapudi et al., 2021].
@@
The literature review reveals several gaps in SA research, particularly in the context of COVID-19 and social media sentiment classification. While existing studies have explored SA using various ML algorithms and lexicon-based approaches, comprehensive investigations remain lacking across diverse datasets, including those from different geographic regions and languages. Additionally, previous research has primarily focused on individual datasets or specific domains, neglecting SA models’ broader applicability and generalizability. Moreover, studies that directly compare different SA techniques, including DL architectures and ensemble methods, are scarce in identifying the most effective approach across various contexts. The need for interpretability and explainability in SA models is also apparent, with few studies incorporating techniques such as SHAP and LIME for insights into model predictions. These gaps highlight the need for more comprehensive and comparative studies encompassing diverse datasets, languages, and evaluation metrics to advance the field of SA effectively.
3 Methodology
@@
Figure 1: This figure depicts the step-by-step methodological framework proposed for tweet SA. It begins with (a) data collection and extension, followed by (b) data cleaning and preprocessing. Subsequently, (c) sentiment labeling into positive, neutral, and negative categories is performed using the ‘cardiffnlp/twitter-roberta-base-sentiment-latest’ pre-trained transformer, and the dataset is split into training, validation, and test sets. The framework proceeds with (d) model development, (e) model benchmarking, and evaluation against baseline models or state-of-the-art approaches. Finally, the process concludes with XAI interpretation techniques applied to gain insights into the model’s predictions.

### v2

@@ Abstract
Sentiment analysis is a pivotal tool in understanding public opinion, consumer behavior, and social trends, underpinning applications ranging from market research to political analysis. However, existing sentiment analysis models frequently encounter challenges related to linguistic diversity, model generalizability, explainability, and limited availability of labeled datasets. To address these shortcomings, we propose the Transformer-based Attention Bidirectional LSTM for Sentiment Analysis (TRABSA) model, a novel hybrid sentiment analysis framework that integrates transformer-based architectures, attention mechanisms, and BiLSTM networks. The TRABSA model leverages the powerful RoBERTa-based transformer model for initial feature extraction, capturing complex linguistic nuances from a vast corpus of tweets. This is followed by an attention mechanism that highlights the most informative parts of the text, enhancing the model’s focus on critical sentiment-bearing elements. Finally, the BiLSTM networks process these refined features, capturing temporal dependencies and improving the overall sentiment classification into positive, neutral, and negative classes. Leveraging the latest RoBERTa-based transformer model trained on a vast corpus of 124M tweets, our research bridges existing gaps in sentiment analysis benchmarks, ensuring state-of-the-art accuracy and relevance. Furthermore, we contribute to data diversity by augmenting existing datasets with 411,885 tweets from 32 English-speaking countries and 7,500 tweets from various US states. This study also compares six word-embedding techniques, identifying the most robust preprocessing and embedding methodologies crucial for accurate sentiment analysis and model performance. We meticulously label tweets into positive, neutral, and negative classes using three distinct lexicon-based approaches and select the best one, ensuring optimal sentiment analysis outcomes and model efficacy. Here, we demonstrate that the TRABSA model outperforms the current seven traditional machine learning models, four stacking models, and four hybrid deep learning models, yielding notable gain in accuracy (94%) and effectiveness with a macro average precision of 94%, recall of 93%, and F1-score of 94%. Our further evaluation involves two extended and four external datasets, demonstrating the model’s consistent superiority, robustness, and generalizability across diverse contexts and datasets. Finally, by conducting a thorough study with SHAP and LIME explainable visualization approaches, we offer insights into the interpretability of the TRABSA model, improving comprehension and confidence in the model’s predictions. Our study results make it easier to analyze how citizens respond to resources and events during pandemics since they are integrated into a decision-support system. Applications of this system provide essential assistance for efficient pandemic management, such as resource planning, crowd control, policy formation, vaccination tactics, and quick reaction programs.
@@
Male domestic violence, Machine learning, Exploratory data analysis, Categorical data, XAI, SHAP, LIME
Introduction
@@
Due to the growth of textual data on social media platforms, news stories, reviews, and consumer feedback, sentiment analysis (SA), a crucial aspect of natural language processing (NLP), has seen growing attention and usage across several domains [1]. Institutes may get crucial insights into public opinion, consumer preferences, market trends, and brand impression by identifying and analyzing feelings conveyed in text [2]. Thus, SA is critical in directing marketing initiatives, product development, company strategies, and reputation management [3]. Additionally, SA is useful in various domains, including politics, healthcare, economics, and the social sciences, where decision-making and policy development depend on a knowledge of human emotions and attitudes.
@@
Despite its broad use, SA still has issues that need more study and creativity. The lack of generalizability and robustness of SA models is one of the primary issues, especially when applying them to various languages, domains, and datasets [4]. Because existing models frequently display different performance levels based on the properties of the data, they are less dependable in real-world situations where the data distribution may change greatly [5]. Furthermore, SA models’ interpretability is still a major worry, particularly in high-stakes scenarios when model predictions are used to make judgments [6]. Deep learning (DL) models are black-box in nature, which makes it difficult to grasp how these models get to their conclusions. This makes it difficult to implement, impedes trust, and holds people accountable for important decision-making processes.
@@
Inspired by these difficulties, this study aims to develop a strong, broadly applicable, and easily interpreted model of SA that will overcome the shortcomings of current techniques. Through improvements in DL, attention mechanisms, and interpretability methodologies, our goal is to develop a model that performs well on various datasets and offers insights into how it makes decisions. This research advances the area of SA by bridging the gap between model performance, interpretability, and practical application. We aim to improve the trustworthiness, transparency, and usefulness of SA models by employing empirical assessments and interpretability studies. This will enable enterprises to make well-informed decisions by relying on dependable sentiment insights.
@@
Our research aims to fill several critical gaps in the existing literature. First, although SA has received a lot of attention—especially regarding social media data—more robust and interpretable models are still required to classify sentiments across various languages and domains accurately. Many current methods are not transparent, scalable, or generalizable, making it difficult to use them in real-world situations. Furthermore, a notable deficiency in current datasets for SA is the absence of representation for various English language usage patterns. Variations in vocabulary, grammar, and contextual usage of English across national boundaries result in subtle discrepancies in the presentation of distinct emotions. This variability presents a problem for SA models, making it difficult to assess sentiments appropriately in various language circumstances. More advanced methods are required to capture minute semantic subtleties and adjust to changing contextual signals since current models may not comprehend sarcasm, context-dependent sentiment changes, or nuanced sentiment expressions.
@@
This study addresses the need for robust and generalizable SA models by proposing the “Transformer-based Attention Bidirectional LSTM for Sentiment Analysis (TRABSA)" model. The TRABSA model integrates the strengths of transformer-based architectures with attention mechanisms and bidirectional long short-term memory (BiLSTM) networks to enhance the performance and adaptability of SA tasks. Our method seeks to capture the broad diversity of English language usage and offer a more thorough knowledge of sentiment expression in various linguistic situations by combining data from several locations into a single dataset. By using this method, TRABSA can more effectively adjust to the subtle differences in English language usage across various groups, which improves the precision and significance of SA findings. We test the performance of the TRABSA model on a range of DL architectures and datasets, including extensive Twitter and external social media datasets, with a particular emphasis on scalability, accuracy, and consistency. Additionally, we do interpretability assessments utilizing the SHAP and LIME approaches to understand the model’s decision-making mechanism. We show the TRABSA model’s generalizability and robustness through our thorough examination, providing a viable method for SA in various real-world situations.
Our research makes eight-fold key contributions:
@@
    We propose the TRABSA model, a novel hybrid sentiment analysis framework that combines transformer-based architectures, attention mechanisms, and BiLSTM networks to improve sentiment analysis performance.
@@
    This research leverages the latest RoBERTa-based transformer model, trained on a vast corpus of 124M tweets, to bridge existing gaps in sentiment analysis benchmarks, ensuring state-of-the-art accuracy and relevance.
@@
    We extended the existing dataset by scraping 411,885 tweets from 32 English-speaking countries to include diversity in the Global Twitter COVID-19 Dataset, acknowledging the varied perspectives and discourse across regions. We scraped an additional 7,500 tweets from different states of the USA to deepen geographical representation in the USA Twitter COVID-19 Dataset, allowing for localized insights and analysis.
@@
    This article thoroughly compares word embedding techniques, establishing the most robust preprocessing and embedding methodologies essential for accurate sentiment analysis and model performance.
@@
    We methodically label tweets using three distinct lexicon-based approaches and rigorously select the most effective one, ensuring optimal sentiment analysis outcomes and model efficacy.
@@
    We conduct extensive experiments to assess the TRABSA model’s performance on the UK Twitter Dataset, benchmarking against 7 traditional machine learning (ML) models, 4 stacking models, and 4 DL models, demonstrating its superiority and versatility.
@@
The rest of this article is structured as follows: Section “Related Works" reviews state-of-the-art literature in SA using ML-DL and interpretability techniques. In section “Methodology", the data collection and preprocessing techniques, unsupervised text labeling, implemented ML and DL models for benchmarking, architecture and methodology of the proposed TRABSA model are described. The section “Results" presents the experimental setup, evaluation metrics, results, and robustness analysis of the TRABSA model. The SHAP and LIME analysis conducted on the TRABSA model are covered in section “Interpretability Analysis". Section “Disucssions" addresses the findings and implications of our investigation. In conclusion, the article is summarized, and future research directions are outlined in section “Conclusions and Future Directions".
Related Works
@@
When it comes to classifying data into positive, neutral, and negative sentiment polarity, SA is essential. Exploring a wide range of emotions is the focus of the emerging domains of SA [7]. Sentiments can be further classified into categories like satisfaction and rage within certain settings, such as political disputes [8]. The development of SA approaches with ambivalence management has allowed classifying emotions into distinct classes, including sorrow, anger, anxiety, excitement, and happiness, leading to more nuanced outcomes [9]. While SA has typically focused on textual data, it has expanded to include multimodal SA, which explains data from devices that employ audio- or audio-visual formats [10]. The extension of SA into multimodal analysis highlights its variety and complexity, creating opportunities for a wide range of NLP applications. The variety of options is further highlighted by the fast growth of NLP, fueled by research in neural networks [11]. Notably, the development of Neurosymbolic AI, which combines symbolic reasoning and deep learning, offers a viable method of improving NLP capabilities [12], highlighting the various paths NLP research is taking. Lexicon-based methods, ML-based methods, and hybrid techniques are the three main methodologies for solving text categorization and emotion detection challenges. Word polarity is used by lexicon-based approaches, and ML techniques see text analysis as a classification problem that may be further divided into supervised, semi-supervised, and unsupervised learning approaches [13]. SA results are frequently improved in real-world applications by combining ML with lexicon-based techniques.
@@
In Ahmed & Ahmed’s work, positive and negative emotions were used to classify gathered fake newspapers using a variety of approaches, including TF-IDF, random forest (RF), Naïve Bayes (NB), etc. [14]. According to their results, out of all the classifiers used, the Naïve Bayes classifier had the best accuracy (89.30%). To identify feelings in the Twitter sentiment 140 datasets, Gaur et al. used TF-IDF feature extraction and the Naïve Bayes Classifier [15]. The model produced improved accuracy (84.44%) and precision when measured using several performance criteria, such as accuracy, recall, and precision. The COVID-19-related data that Qi & Shabrina (2023) examined came from Twitter users in major English cities [16]. They conducted a comparative analysis of ML models, including Vader and Textblob, RF, support vector classification (SVC), and multinominal Naïve Bayes (MNB) models. According to the results of their investigation, SVC with TF-IDF demonstrated better accuracy than the other models. To assess opinions about Saudi cruises, Al Sari et al. created three different datasets from social media platforms [17]. With oversampled Snapchat data, they used ML techniques, including RF, MLP, NB, voting, SVM, and the n-grams feature extraction approach to reach 100% accuracy with the RF algorithm. A customized approach for explicit negation detection was presented by Mukherjee et al. [18]. They used TF-IDF for feature extraction and various ML techniques, including NB, SVM, and Artificial Neural Networks (ANN), to analyze sentiment in Amazon reviews. According to their research, ANNs using negative classifiers had the best accuracy (96.32%). Using reviews from an international hotel, Noori developed a unique algorithm for classifying client sentiment [19]. Following the processing of the reviews, document vectors were created using the TF-IDF extractor and trained using SVM, ANN, NB, k-nearest neighbor (KNN), decision tree (DT), and C4.5 models. Outperforming other models, the DT model scored the highest accuracy (98.9%) with 1800 features. Using N-gram extraction, Zahoor & Rohilla compared NB, SVM, RF, and long short-term memory networks (LSTM) classifiers on preprocessed datasets [20]. In most datasets, including the BJP and ML Khattar datasets, NB showed the best accuracy. To turn COVID-19-related tweets into a text corpus and determine the most common terms using N-grams, Samuel et al. used logistic regression (LR) and NB models [21]. Their results showed that for short tweets, NB and LR had peak accuracy rates of 91% and 74%, respectively. For lengthier tweets, both models performed pretty poorly. Using Maximum Entropy (ME), SVM, and LSTM models, Kumar et al. examined the effects of age and gender on customer reviews [22]. LSTM used word2vec, but the NB, ME, and SVM algorithms used Bag of Words (BOW) feature extraction. For female data, the over-50 age group showed the highest accuracy. SVM and MNB with TF-IDF extraction were used by Zarisfi Kermani et al. on four Twitter datasets, and they proposed semantic scoring techniques to represent features in the vector space [23]. According to their findings, the suggested technique outperformed the MNB algorithm in three datasets, with the STS dataset showing the greatest MNB performance.
@@
In 13 languages with different Indic scripts, Bansal et al. looked at the identification of objectionable language [24]. They assessed four sophisticated transformer-based models and contrasted the Transformer-based method with traditional ML models. Out of all of them, XLM-RoBERTa with BiGRU performed better. Furthermore, adding emoji embeddings to XLM-RoBERTa improved the model’s efficacy even further. Due to the combined dataset’s code-mixing, training using datasets from 13 Indic languages performed better than training with separate models. Gupta et al. [25] presented a unique emotion analysis approach for real-time COVID-19 tweets, examining eight emotions in different domains. The analysis of tweets from India showed changes in emotional reactions, such as less happiness except for nature. Because of their commitment, teachers’ faith in education has grown. In terms of precision and recall, the method by Gupta et al. produced aspect-based graphical and textual summaries from mobile reviews [26], outperforming baseline approaches. Using Twitter data from the Delhi Election 2020, Gupta et al. conducted political echo chamber experiments and investigated the elements that contribute to the creation of echo chambers as well as the role played by users of opposing parties in promoting partisan material [27]. Gupta and colleagues employed ML algorithms and lexicon-based methods to assess sentiment in Hindi tweets. They found that an integrated CNN-RNN-LSTM model produced an accuracy of 85% [28]. Basiri et al. investigated attitudes about the epidemic in eight different nations using DL classification algorithms, and their findings showed distinct sentiment patterns and relationships with pandemic indicators [29]. Using the BERT model, Hayawi et al. achieved excellent accuracy in their ML-based method for spotting COVID-19 vaccination disinformation [30]. Using BERT, Vishwamitra et al. were able to detect hate speech connected to elderly individuals and the Asian community on Twitter during the pandemic [31]. They were able to identify separate word connections for various hate speech datasets. Before and after the initial COVID-19 case announcement, Chen et al. monitored conversations in Luxembourg about policy and daily life; post-announcement, travel-related issues dominated, perhaps because of the region’s large immigrant population [32]. To emphasize changing emotions over time, Kabir et al. used ML for word extraction and emotion categorization in COVID-19 tweets [33]. To categorize COVID-19-related Twitter postings, Valdes et al. created a BERT-based model [34], proving the use of domain-specific data for improved performance. During the pandemic, Tziafas et al. used an ensemble architecture to recognize false information, using transformer-based encoders to achieve high accuracy [35]. Sadia et al. obtained high assessment scores using BERT to conduct SA of COVID-19 tweets [36]. Song et al. analyzed several facets of misinformation diffusion and created a model to categorize misinformation related to COVID-19 [37]. Using NLP models, Hossain et al. assessed a dataset for COVID-19-related misinformation detection, offering preliminary benchmarks for advancement [38]. During the COVID-19 lockdown, Chintalapudi et al. used BERT to assess sentiment in Indian tweets, and they showed better accuracy than other models [39].
@@
The literature review reveals several gaps in SA research, particularly in the context of COVID-19 and social media sentiment classification. While existing studies have explored SA using various ML algorithms and lexicon-based approaches, comprehensive investigations remain lacking across diverse datasets, including those from different geographic regions and languages. Additionally, previous research has focused on individual datasets or specific domains, neglecting SA models’ broader applicability and generalizability. Moreover, studies that directly compare different SA techniques, including DL architectures and ensemble methods, are scarce in identifying the most effective approach across various contexts. The need for interpretability and explainability in SA models is also apparent, with few studies incorporating techniques such as SHAP and LIME for insights into model predictions. These gaps highlight the need for more comprehensive and comparative studies encompassing diverse datasets, languages, and evaluation metrics to advance the SA field effectively.
Methodology
@@
Figure 1: This figure depicts the step-by-step methodological framework proposed for tweet sentiment analysis. It begins with (a) data collection and extension, followed by (b) data cleaning and preprocessing. Subsequently, (c) sentiment labeling into positive, neutral, and negative categories is performed using the ‘cardiffnlp/twitter-roberta-base-sentiment-latest’ pre-trained transformer, and the dataset is split into training, validation, and test sets. The framework proceeds with (d) model development, (e) model benchmarking, and evaluation against baseline models or state-of-the-art approaches. Finally, the process concludes with XAI interpretation techniques applied to gain insights into the model’s predictions.